### v1

@@ 4.1 Setting Up
We randomly sample 2.5k instructions from WildChat (Zhao et al., ) to collect stepwise preference pairs, which we then combine with Ultrafeedback (Cui et al., 2024) for DPO training. The learning rate is set to 1e-6, and we employ a cosine learning rate scheduler. The maximum sequence length is configured to 32,768 through packing, with a random seed set to 42, and training is conducted for 250 steps.
@@
4.2 Main Results
Table 2: Performance comparison across more long-form and general benchmarks. LongGenBench can be used to evaluate output lengths up to 32k. The other benchmarks assess the model’s performance on more general tasks like helpfulness and reasoning. For TruthfulQA, we report partition “MC1" and “MC2". For each task, all three methods use the same decoding settings, and we have set a random seed to ensure reproducibility.
@@ 4.2 Main Results
Table 3: Ablation on refinement methods and “w/o critique" stands for without critiques meaning MCTS is applied alone. “Self-critique" refers to critiques generated by the model itself. To verify generalization, we set different values of $\eta$ and report the average result.
@@ 4.4 Comparision with Different Critic Methods
To enable a more thorough comparison, we set multiple values for $\eta$ in Equation 6. Specifically, we set $\eta$ to {2.0, 2.5, 3.0} and report the average performance in Table 3. We detailed the results in Table 7 and 8. Self-critique exhibits performance fluctuations which may be because the generator’s internal knowledge is insufficient, making it difficult to distinguish high-quality steps in long-form generation.
@@ 5.1 Reliability of Evaluation
Human Evaluation In addition to utilizing GPT-4o, we conduct a human evaluation to assess the generated text in terms of diversity, consistency, and informative detailed guidelines can be seen in A.4. We compare the responses generated by LongWriter-Llama and LongWriter-Qwen with those produced by the same models trained using LongDPO. Three independent annotators, who are undergraduate and graduate students, are tasked with comparing the response pairs and evaluating them as win, tie, or lose. The results, present in Table 4, indicate that our responses are rated as superior by the human judges. Additionally, Table 5 shows the agreement among the three judges, demonstrating a high level of consistency in their evaluations.
5.2 Case Study
We analyze the attention distribution across models and observe that, in week 15, LongWriter-Llama fails to attend to “farmers market." However, after applying LongDPO, it successfully does so. We find that a small number of attention heads have attended to “farmers market," with over 1% of attention heads scoring above 0.5. However, the LongWriter model does not exhibit a similar pattern. This behavior may be linked to retrieval heads (Wu et al., 2024b). We also provide examples in Figure 7 and 8 to show factual correctness after applying LongDPO.
@@ 6 Conclusion
In this paper, we propose LongDPO to enhance long-form generation abilities for LLMs. First, we construct stepwise preference pairs by leveraging process supervision with MCTS, incorporating a global memory pool to maintain factual consistency, and utilizing external critiques to refine the chosen candidates which are of low-rewards in the collected preference pairs. Then, LongDPO utilizes step-level preference data for DPO training. This approach enhances performance in long-form generation tasks while maintaining near-lossless performance on several general tasks. Furthermore, we conduct an in-depth analysis to show models trained by LongDPO can generate responses that align more closely with humans.
Impact Statement
This work proposes methods for generating high-quality long-form text, which may help make human content creation easier. We have introduced approaches aimed at improving the factual accuracy of the generated content. We believe that long-form generation is a promising direction worth exploring which may have an impact on legal, code content generation.
Appendix A Appendix
@@ A.1 Reward Evaluation Templates
Reward Evaluation Template You are an expert at evaluating the quality of text. As an impartial evaluator, please assess the assistant’s response to a user’s requirements. Now, you will receive specific principles that provide the criteria for evaluating the response. Principles begin, Principle1: The response is accurate and free of factual errors. Principle2: The response meets the user’s purpose and needs. Principle3: The response is non-toxic and safe. Principle4: The response meets the user’s formatting requirements and maintains logical consistency. Principle5: The response contains diverse and comprehensive information with minimal repetition. Principle6: The response provides an excellent reading experience. Principle7: The response is insightful and provides the user with additional avenues for thought. Principles end. In the next, you will receive detailed guidelines to help you rate the response according to each principle. Now, guidelines begin 5: A perfect response with no improvement needed. The content is comprehensive, accurate, clear, and well-structured. The response fully addresses all aspects of the question or need without any omissions or errors. 4: A very good response with minor issues. It is almost perfect but may have slight areas that could be improved, such as minor details that are unclear or a small omission. Overall, it still meets the need effectively. 3: An acceptable response that generally meets the question or need but has noticeable shortcomings. The content might be incomplete or unclear, or there may be minor grammar or logical errors. It needs improvement but is still functional. 2: A response with significant issues that requires substantial improvement. The content is incomplete, unclear, or contains major errors, omissions, or misunderstandings. It does not fully satisfy the request. 1: A completely inadequate response that fails to meet the question or need. It contains serious errors or misunderstandings and cannot provide useful help. Guidelines end. Now, you will receive the user request and the assistant’s response to evaluate. <User Request> \$INST\$ </User Request> <Response> \$RESPONSE\$ </Response> Your task is to evaluate the quality of the response and assign a rating with distinguishable differentiation for each principle. When rating, please carefully read the guidelines and ensure your ratings fully adhere to them. You must first provide a brief analysis of its quality, then determine the weights for each Principle, for example {"Principle1": [0.2,0.2,0.2,0.2,0.2]} represents the final score is 0.2 * 1 + 0.2 * 2 + 0.2 * 3 + 0.2 * 4 + 0.2 * 5 = 3. The output must strictly follow the JSON format: {"Analysis":..., "Principle1":[..,..,..,..,..], "Principle2":[..,..,..,..,..], "Principle3":[..,..,..,..,..], "Principle4":[..,..,..,..,..], "Principle5":[..,..,..,..,..], "Principle6":[..,..,..,..,..], "Principle7":[..,..,..,..,..]}. You do not need to consider whether the response meets the user’s length requirements in your evaluation. Ensure that only one integer or float is output for each principle.
@@ A.4 Guidelines for Human Annotation
A.5 More Evaluation Results
@@
Table 7: Results on changing $\eta$ using llama-based backbones
@@
Table 8: Results on changing $\eta$ using Qwen-based backbones

### v2

@@ 4.1 Setting Up
We randomly sample 2.5K instructions from WildChat (Zhao et al., ) to collect stepwise preference pairs, which we then combine with UltraFeedback (Cui et al., 2024) for training. For data from UltraFeedback, we use vanilla DPO. The learning rate is set to 1e-6, with a cosine learning rate scheduler. The maximum sequence length is 32,768 through packing, with a random seed set to 42, and training for 250 steps. We use Xtuner  [^6] https://github.com/InternLM/xtuner for training.
@@
Table 2: Performance comparison across more long-form and general benchmarks. LongGenBench can be used to evaluate output lengths up to 32k. For TruthfulQA, we report partition “MC1” and “MC2”. For each task, all three methods use the same decoding settings, and we have set a random seed to ensure reproducibility.
@@
Table 3: Ablation on refinement methods and “w/o critique” stands for without critiques meaning MCTS is applied alone. “Self-critique” refers to critiques generated by the model itself. To verify generalization, we set different values of $\eta$ and report the average result.
4.2 Main Results
In addition to the 7B-sized model, we also conducted experiments on larger models and compared them with more advanced open-source models. Detailed results can be seen in Table 13.
@@ 4.3 Generalization on more long-form and general benchmarks
LLM-AggreFact (without threshold tuning)
Model Name AggreFact TofuEval Wice Reveal Claim Verify Fact Check Expert QA Lfqa RT Avg
CNN XSum MediaS MeetB
   LongWriter-Qwen 52.71 71.55 73.33 75.83 74.40 87.73 70.18 74.61 60.56 84.61 76.65 72.92
   w/o Memory 52.03 69.31 72.16 75.38 76.07 87.58 68.46 74.94 60.27 83.36 75.70 72.30
   w/ Memory 54.36 73.20 73.28 76.25 74.92 88.31 70.87 73.79 61.23 86.76 77.39 73.67
Table 4: Performance (BAcc) of evaluator models on the test split of LLM-AggreFact. “RT” represents RAGTruth.
@@ 4.4 Comparision with Different Critic Methods
To enable a more thorough comparison, we set multiple values for $\eta$ in Equation 6. Specifically, we set $\eta$ to {2.0, 2.5, 3.0} and report the average performance in Table 3. We detailed the results in Table 9 and 10. Self-critique exhibits performance fluctuations which may be because the generator’s internal knowledge is insufficient, making it difficult to distinguish high-quality steps.
4.5 Effects of the Memory Pool
We assess the effectiveness of the memory pool using the LLM-AggreFact (Tang et al., 2024), which includes a variety of fact-checking tasks. The results are presented in Table 4. Without using memory to collect data and training directly, the fact-checking scores decreased. However, after incorporating memory, the model’s fact-checking ability improved.
Models LongGenBench
CR STC1 STC2
LongWriter-Llama
    w/o Stepwise 67.89 25.36 17.29
    w/ Stepwise 69.38 27.59 18.45
LongWriter-Qwen
    w/o Stepwise 97.42 31.95 31.44
    w/ Stepwise 98.51 33.07 32.52
Table 5: Performance comparison in LongGenBench.
4.6 Effects of Stepwise Learning
We evaluate the impact of stepwise learning on long-form generation using LongGenbench. The results are shown in Table 5. We use the same training data. The difference between the methods is that “w/o Stepwise” refers to training with vanilla DPO, while “w/ Stepwise” refers to training with the LongDPO objective. Stepwise learning is beneficial for learning long-form generation. The detailed results shown in Table 11.
@@ 5.1 Reliability of Evaluation
Human Evaluation In addition to utilizing GPT-4o, we conduct a human evaluation to assess the generated text in terms of diversity, consistency, and informative detailed guidelines can be seen in A.4. We compare the responses generated by LongWriter-Llama and LongWriter-Qwen with those produced by the same models trained using LongDPO. Three independent annotators, who are undergraduate and graduate students, are tasked with comparing the response pairs and evaluating them as win, tie, or lose. The student participants all possess a bachelor’s or master’s degree and are from top universities and have two years of experience in NLP. The results, present in Table 6, indicate that our responses are rated as superior by the human judges. Additionally, Table 7 shows the agreement among the three judges, demonstrating a high level of consistency in their evaluations.
5.2 Case Study
We analyze the attention distribution across models and observe that, in week 15, LongWriter-Llama fails to attend to “farmers market.” However, after applying LongDPO, it successfully does so. We find that a small number of attention heads have attended to “farmers market,” with over 1% of attention heads scoring above 0.5. However, the LongWriter model does not exhibit a similar pattern. This behavior may be linked to retrieval heads (Wu et al., 2024b). We also provide examples in Figure 7 and 8 to show factual correctness after applying LongDPO.
6 Disscussion
LongDPO focuses on long-form tasks (e.g., Creative Writing), which, unlike tasks such as math and coding, do not have a ground truth. It is more challenging to assess the reward precisely. Different from existing literature in reinforcement learning, which can rely on rule-based rewards or process reward models, we take into full consideration the characteristics of natural language and have carefully designed seven principles for evaluating the reward.
@@ 7 Conclusion
In this paper, we propose LongDPO which incorporate process supervision with MCTS to collect better preference pairs with a memory pool to maintain factual consistency and leverages external critiques to refine low-quality candidates in long-form generation. LongDPO enhances performance in long-form generation tasks (e.g. LongBench-Write) while maintaining near-lossless performance on several general tasks.
Limitations
We have validated the effectiveness of LongDPO in generating text of 32K length. However, due to the limitations of current benchmarks, it is challenging to evaluate longer generation lengths. In the future, we plan to test the performance of LongDPO further on longer benchmarks.
Acknowledgements
This work was supported by the National Science and Technology Major Project (No. 2022ZD0117800).
Appendix A Templates and Guidelines
@@ A.1 Reward Evaluation Templates
Reward Evaluation Template You are an expert at evaluating the quality of text. As an impartial evaluator, please assess the assistant’s response to a user’s requirements. Now, you will receive specific principles that provide the criteria for evaluating the response. Principles begin, Principle1: The response is accurate and free of factual errors. Principle2: The response meets the user’s purpose and needs. Principle3: The response is non-toxic and safe. Principle4: The response meets the user’s formatting requirements and maintains logical consistency. Principle5: The response contains diverse and comprehensive information with minimal repetition. Principle6: The response provides an excellent reading experience. Principle7: The response is insightful and provides the user with additional avenues for thought. Principles end. In the next, you will receive detailed guidelines to help you rate the response according to each principle. Now, guidelines begin 5: A perfect response with no improvement needed. The content is comprehensive, accurate, clear, and well-structured. The response fully addresses all aspects of the question or need without any omissions or errors. 4: A very good response with minor issues. It is almost perfect but may have slight areas that could be improved, such as minor details that are unclear or a small omission. Overall, it still meets the need effectively. 3: An acceptable response that generally meets the question or need but has noticeable shortcomings. The content might be incomplete or unclear, or there may be minor grammar or logical errors. It needs improvement but is still functional. 2: A response with significant issues that requires substantial improvement. The content is incomplete, unclear, or contains major errors, omissions, or misunderstandings. It does not fully satisfy the request. 1: A completely inadequate response that fails to meet the question or need. It contains serious errors or misunderstandings and cannot provide useful help. Guidelines end. Now, you will receive the user request and the assistant’s response to evaluate. <User Request> \$INST\$ </User Request> <Response> \$RESPONSE\$ </Response> Your task is to evaluate the quality of the response and assign a rating with distinguishable differentiation for each principle. When rating, please carefully read the guidelines and ensure your ratings fully adhere to them. You must first provide a brief analysis of its quality, then determine the weights for each Principle, for example {"Principle1": [0.2,0.2,0.2,0.2,0.2]} represents the final score is 0.2 * 1 + 0.2 * 2 + 0.2 * 3 + 0.2 * 4 + 0.2 * 5 = 3. The output must strictly follow the JSON format: "Analysis":…, "Principle1":[..,..,..,..,..], "Principle2":[..,..,..,..,..], "Principle3":[..,..,..,..,..], "Principle4":[..,..,..,..,..], "Principle5":[..,..,..,..,..], "Principle6":[..,..,..,..,..], "Principle7":[..,..,..,..,..]. You do not need to consider whether the response meets the user’s length requirements in your evaluation. Ensure that only one integer or float is output for each principle.
@@ A.4 Guidelines for Human Annotation
Appendix B More Evaluation Results
@@
Table 9: Results on changing $\eta$ using llama-based backbones
@@
Table 10: Results on changing $\eta$ using Qwen-based backbones
Models LongGenBench (16K) LongGenBench (32K)
CR STC1 STC2 CR STC1 STC2
LongWriter-Llama
    w/o Stepwise 67.89 25.36 17.29 67.79 31.85 21.67
    w/ Stepwise 69.38 27.59 18.45 68.35 33.69 22.15
LongWriter-Qwen
    w/o Stepwise 97.42 31.95 31.44 83.78 28.82 23.24
    w/ Stepwise 98.51 33.07 32.52 84.95 29.86 24.32
Table 11: Performance comparison in LongGenBench.
@@
Models [0, 500) [500, 2k) [2k, 4k) [4k, 20k) Average
$S_{l}$ $S_{q}$ $S_{l}$ $S_{q}$ $S_{l}$ $S_{q}$ $S_{l}$ $S_{q}$ $S_{l}$ $S_{q}$
Llama3.1-8B-instruct 89.70 84.60 78.20 80.60 29.20 76.10 0 57.60 56.80 76.30
Llama3.1-70B-instruct 90.80 84.80 88.60 84.40 14.90 84.50 0 78.00 59.00 80.30
GPT-4o 92.10 93.10 92.20 93.50 53.00 92.80 6.20 81.20 67.80 90.90
QWQ 89.10 94.58 90.59 94.31 33.13 93.75 0.26 89.39 53.27 93.01
Qwen-2.5-14B 88.39 89.77 81.83 91.91 71.38 87.50 19.57 82.35 65.29 87.88
  w/ DPO 91.72 90.53 81.79 92.72 68.94 86.50 18.33 80.20 65.19 87.48
  w/ LongDPO 91.75 91.25 85.69 90.53 78.79 89.01 21.50 86.04 69.43 89.21
Table 13: More evaluation results of larger models on LongBench-Write-en.